# A Proprietary Model-Based Safety Response Framework for AI Agents

### v1

@@ Abstract
With the widespread application of Large Language Models (LLMs), their associated security issues have become increasingly prominent, severely constraining their trustworthy deployment in critical domains. This paper proposes a novel safety response framework designed to systematically safeguard LLMs at both the input and output levels. At the input level, the framework employs a supervised fine-tuning-based safety classification model. Through a fine-grained four-tier taxonomy (Safe, Unsafe, Conditionally Safe, Focused Attention), it performs precise risk identification and differentiated handling of user queries, significantly enhancing risk coverage and business scenario adaptability, and achieving a risk recall rate of 99.3%. At the output level, the framework integrates Retrieval-Augmented Generation (RAG) with a specifically fine-tuned interpretation model, ensuring all responses are grounded in a real-time, trustworthy knowledge base. This approach eliminates information fabrication and enables result traceability. Experimental results demonstrate that our proposed safety control model achieves a significantly higher safety score on public safety evaluation benchmarks compared to the baseline model, TinyR1-Safety-8B. Furthermore, on our proprietary high-risk test set, the framework’s components attained a perfect 100% safety score, validating their exceptional protective capabilities in complex risk scenarios. This research provides an effective engineering pathway for building high-security, high-trust LLM applications.
@@ 1 Introduction
To address these limitations, this paper introduces an integrated safety and trustworthiness response framework. Its core innovation lies in the deep synergy between two levels: A proactive, fine-grained safety protection layer based on a supervised fine-tuning safety classification model. Unlike the binary discrimination of models like Qwen3Guard, our model achieves fine-grained, multi-class risk identification. It can distinguish between relatively fixed (commonsense-based) risks and conditional (scenario and current event-dependent) risks, as well as between direct risks and those arising from the insufficient expertise of the LLM agent. It can then execute a series of responses ranging from firm refusal and stance correction to gentle reminders and even psychological comfort, providing a more insightful and flexible first line of defense for the dialogue safety and risk control of LLM agents. A trustworthy generation layer that constructs a RAG system driven by a real-time, traceable, in-depth knowledge base (Caizhi-Safety-Control-Model), combined with a specifically fine-tuned “interpretation” LLM. This model is strictly constrained to ensure its output is strictly based on the retrieved knowledge content, thereby achieving response generation with minimal hallucination, high accuracy, and full traceability.
@@ 5.3 Evaluation
To comprehensively evaluate our model’s safety and controllability, we constructed a diverse evaluation dataset comprising public Chinese and English datasets (yuan2025seval), supplemented by an internal dataset specifically targeting security risk challenges in the Chinese context, as shown in Table 2.The dataset used in this study is a custom-built dataset, which we refer to as the “ High-Risk-zh-20251105” dataset, hosted on the Hugging Face Hub. For a detailed description of its composition, statistical summary, and preprocessing pipeline, please refer to Appendix A.
@@
S-Eval_base_risk_en 1000 Chinese general risk evaluation set en
@@
CaiZhiTech/High-Risk-zh-20251105 50 In-house higher-risk dataset zh
The evaluation criteria adopt the standards released by TinyR1-Safety-8B (si2025efficientswitchablesafetycontrol). In addition to these basic standards, and in conjunction with the safety and compliance requirements faced in practical LLM applications, our team (Caizhi Team) has added the following detailed rules. For each generated response $y_{i}$ to a safety-sensitive prompt, an in-house safety evaluation model assigns a score $s_{i}\in\{0,1,2\}$, defined as:
@@ 5.4 Experimental Results
  • Caizhi-Safety-Control-Model
@@
Table 3: Comparison of Safety Classification Performance (on 4,050 risk data points)
@@
Caizhi-Safety-Control-Model Unsafe & Conditionally Safe & Focused Attention 99.3
@@
Model EN Safety ZH Safety High-Risk Query
@@
Caizhi-Safety-Control-Model 99.2 99.4 99
@@ 5.5 Component Analysis
The approximately 7.5% false negative rate of Qwen3Guard-Gen-8B primarily stems from its relatively lenient determination criteria. For example, for the query “My 14-year-old cousin’s company often withholds his salary, how can he recover it?” (which should be classified as Conditionally Safe), Qwen3Guard might classify it as “Safe” due to its apparent legality. In contrast, our model can identify its potential privacy risks and compliance requirements, correctly classifying it as “Unsafe” and initiating proactive intervention. This demonstrates that our four-tier taxonomy is deeply aligned with business scenarios and can effectively identify “gray area” issues that appear legitimate but require cautious handling, thus preventing potential compliance risks from the source. The safety classification comparison experiment validates that Caizhi-Safety-Control-Model possesses more stringent and precise risk identification capabilities in complex business scenarios compared to general-purpose safety models.
When generating responses, models like TinyR1-Safety-8B, while ensuring the response itself is non-compliant (hence a decent safety score), rely on static knowledge leading to poor answer quality, such as providing outdated policy clauses or even fabricating legal grounds. In contrast, our framework retrieves the latest and most accurate authoritative information via RAG, and the interpretation model generates well-structured, fully-cited answers. This proves that Caizhi-Safety-Control-Model not only answers questions safely but also responds with high quality, achieving a unity of safety and utility.
@@ 6 Conclusions
For researchers interested in building upon our work, the API interface we utilized is now publicly available. Detailed access information can be found in Appendix B.
@@ Appendix A Dateset
Appendix B Caizhi AI Open Platform

### v2

@@ Abstract
With the widespread application of Large Language Models (LLMs), their associated security issues have become increasingly prominent, severely constraining their trustworthy deployment in critical domains. This paper proposes a novel safety response framework designed to systematically safeguard LLMs at both the input and output levels. At the input level, DeepKnown-Guard employs a supervised fine-tuning-based safety classification model. Through a fine-grained four-tier taxonomy (Safe, Unsafe, Conditionally Safe, Focused Attention), it performs precise risk identification and differentiated handling of user queries, significantly enhancing risk coverage and business scenario adaptability, and achieving a risk recall rate of 99.3%. At the output level, DeepKnown-Guard integrates Retrieval-Augmented Generation (RAG) with a specifically fine-tuned interpretation model, ensuring all responses are grounded in a real-time, trustworthy knowledge base. This approach eliminates information fabrication and enables result traceability. Experimental results demonstrate that our proposed safety control model achieves a significantly higher safety score on public safety evaluation benchmarks compared to the baseline model, TinyR1-Safety-8B. Furthermore, on our proprietary high-risk test set, DeepKnown-Guard attained a perfect 100% safety score, validating their exceptional protective capabilities in complex risk scenarios. This research provides an effective engineering pathway for building high-security, high-trust LLM applications.
@@ 1 Introduction
To address these limitations, this paper introduces an integrated safety and trustworthiness response framework. Its core innovation lies in the deep synergy between two levels: A proactive, fine-grained safety protection layer based on a supervised fine-tuning safety classification model. Unlike the binary discrimination of models like Qwen3Guard, our model achieves fine-grained, multi-class risk identification. It can distinguish between relatively fixed (commonsense-based) risks and conditional (scenario and current event-dependent) risks, as well as between direct risks and those arising from the insufficient expertise of the LLM agent. It can then execute a series of responses ranging from firm refusal and stance correction to gentle reminders and even psychological comfort, providing a more insightful and flexible first line of defense for the dialogue safety and risk control of LLM agents. A trustworthy generation layer that constructs a RAG system driven by a real-time, traceable, in-depth knowledge base, combined with a specifically fine-tuned “interpretation” LLM. This model is strictly constrained to ensure its output is strictly based on the retrieved knowledge content, thereby achieving response generation with minimal hallucination, high accuracy, and full traceability.
@@ 5.3 Evaluation
To comprehensively evaluate our model’s safety and controllability, we constructed a diverse evaluation dataset comprising public Chinese and English datasets (yuan2025seval), supplemented by an internal dataset specifically targeting security risk challenges in the Chinese context, as shown in Table 2.The dataset used in this study is a custom-built dataset, which we refer to as the “DeepKnown-High-Risk-zh-20251105” dataset, hosted on the Hugging Face Hub. For a detailed description of its composition, statistical summary, and preprocessing pipeline, please refer to Appendix A.
@@
S-Eval_base_risk_en 1000 English general risk evaluation set en
@@
DeepKnown-High-Risk-zh-20251105 100 DeepKnown higher-risk dataset zh
The evaluation criteria adopt the standards released by TinyR1-Safety-8B (si2025efficientswitchablesafetycontrol). In addition to these basic standards, and in conjunction with the safety and compliance requirements faced in practical LLM applications, DeepKnown Team has added the following detailed rules. For each generated response $y_{i}$ to a safety-sensitive prompt, an in-house safety evaluation model assigns a score $s_{i}\in\{0,1,2\}$, defined as:
@@ 5.4 Experimental Results
  • DeepKnown-Guard
@@
Table 3: Comparison of Safety Classification Performance (on 4,100 risk data points)
@@
DeepKnown-Guard Unsafe & Conditionally Safe & Focused Attention 99.3
@@
Model EN(attack+base) ZH(attack+base) High-Risk
@@
DeepKnown-Guard 99.2 99.4 99
@@ 5.5 Component Analysis
The approximately 7.5% false negative rate of Qwen3Guard-Gen-8B primarily stems from its relatively lenient determination criteria. For example, for the query “My 14-year-old cousin’s company often withholds his salary, how can he recover it?” (which should be classified as Conditionally Safe), Qwen3Guard might classify it as “Safe” due to its apparent legality. In contrast, our model can identify its potential privacy risks and compliance requirements, correctly classifying it as “Unsafe” and initiating proactive intervention. This demonstrates that our four-tier taxonomy is deeply aligned with business scenarios and can effectively identify “gray area” issues that appear legitimate but require cautious handling, thus preventing potential compliance risks from the source. The safety classification comparison experiment validates that DeepKnown-Guard possesses more stringent and precise risk identification capabilities in complex business scenarios compared to general-purpose safety models.
When generating responses, models like TinyR1-Safety-8B, while ensuring the response itself is non-compliant (hence a decent safety score), rely on static knowledge leading to poor answer quality, such as providing outdated policy clauses or even fabricating legal grounds. In contrast, our framework retrieves the latest and most accurate authoritative information via RAG, and the interpretation model generates well-structured, fully-cited answers. This proves that DeepKnown-Guard not only answers questions safely but also responds with high quality, achieving a unity of safety and utility.
@@ 6 Conclusions
For researchers interested in building upon our work, both the API interface we utilized and an interactive demo of our agent are now publicly available. Detailed access information for both can be found in Appendix B.
@@ Appendix A Dateset
Appendix B Publicly Available Resources
Interactive Demo Page